# Spin-chain multichannel Kondo model via image impurity boundary condition

### v1

@@ IV Total Impurity Spin
Another signature of MCK physics in our model is exhibited by the total spin of the impurity and image impurity. We first find that the total spin of the whole system is spin 00 for all $0<J_{K}^{\prime}=J_{K}\leq 1$ and $M=1,2,3$. This singlet formed by the impurity spin and the rest of the spins is similar to the electronic multichannel Kondo model where the impurity spin was connected to entanglement negativity [61]. The negativity was also numerically calculated for the spin-chain 1CK case [65]. We calculate the total spin of the impurity and the image impurity, or equivalently the total spin of the $M\times L$ bulk spins because the whole system has spin 0. Due to the $\text{SU}(2)$ symmetry, it is sufficient to calculate the $z$-component,
⟨$\langle(\mathrm{S}^{z}_{\text{tot}})^{2}\rangle=\langle(\mathrm{S}_{0}^{z}+\mathrm{S}_{0}^{{}^{\prime}z})^{2}\rangle=\frac{1}{2}+2\langle\mathrm{S}_{0}^{z}\mathrm{S}_{0}^{{}^{\prime}z}\rangle\,,$ (9)
where ⟨$\langle\mathrm{S}_{0}^{z}\mathrm{S}_{0}^{z}\rangle=\langle\mathrm{S}_{0}^{{}^{\prime}z}\mathrm{S}_{0}^{{}^{\prime}z}\rangle=1/4$ by $\text{SU}(2)$ symmetry.
@@
We find that the total impurity spin Eq. (9) depends on the system size and Kondo coupling through the ratio $L/\xi_{K}$. The finite-size scaling for ⟨$\langle(\mathrm{S}^{z}_{\text{tot}})^{2}\rangle$ is shown in Fig. 2(b). In the $L/\xi_{K}\to\infty$ limit, ⟨$\langle(\mathrm{S}^{z}_{\text{tot}})^{2}\rangle\to 1/2$, which corresponds to uncorrelated impurity and image impurity. In the bulk, we get exactly the same value for the $z$-component total spins, which suggests that two independent spin-1/2 Kondo clouds are formed in the bulk. Because $\xi_{K}$ is about the size of the Kondo screening cloud, we can think of it the following way: One of the Kondo clouds is for the impurity, the other is for the image-impurity and they are decoupled from each other because their sizes $\xi_{K}$ are very small compared to the chain length $L$. In this limit, we find ⟨$\langle(\mathrm{S}^{z}_{\text{tot}})^{2}\rangle$ to deviate from 1/2 in a Fermi liquid power law $1/(L/\xi_{K})^{2}$ for the $M=1$ case and non-Fermi liquid power laws $1/(L/\xi_{K})$ for $M=2$ and $1/(L/\xi_{K})^{4/5}$ for $M=3$, see Fig. 2(b). These finite-size corrections can be understood to arise in second order perturbation theory in the least irrelevant operator with scaling dimension $2$ for $M=1$ and $1+2/(M+2)$ for $M\geq 2$ [15, 16]. A similar scaling behavior can be found in Ref. [61] for the negativity of the impurity and the bulk electrons. In the opposite limit $L/\xi_{K}\to 0$, the correlation between the impurity and the image-impurity mediated by their Kondo clouds becomes strong. Eventually, the impurity and the image-impurity form a singlet and the $M\times L$ bulk spins form another singlet.
V Effect of Anisotropy
To implement anisotropy, we use the Hamiltonian Eq. 6 but replace ${\bf\it\mathrm{S}}_{i}\cdot{\bf\it\mathrm{S}}_{j}\to\mathrm{S}_{i}^{x}\mathrm{S}_{j}^{x}+\mathrm{S}_{i}^{y}\mathrm{S}_{j}^{y}+\Delta\mathrm{S}_{i}^{z}\mathrm{S}_{j}^{z}$ where $\Delta=1$ gives the isotropic case. We will consider only the critical regime −$-1<\Delta\leq 1$ [52].
@@ V Effect of Anisotropy
Figure 3: The effective ground state degeneracy $g=\exp(S_{\text{imp}})$ versus XXZ anisotropy parameter for $M=1,2,3,4$ and $L=50,50,50,30$ respectively. The power law $g\sim d^{(M-1)/2}$ is observed where the fit lines are for the $\Delta\leq 0$ ($d<1/2$) points (Luttinger liquid parameter $K>1$ [68]). The fit does not work well near $\Delta=1$ ($d=1$), where it yields $g\sim\sqrt{M}$ instead of Eq. (1).
@@
We use DMRG simulations to find $g(M,\Delta)$ for $M=1,\dots,4$, shown in Fig. 3. We find that $g(M,\Delta)\sim d(\Delta)^{\alpha}$ closely follows a power law with the exponent $\alpha\approx(M-1)/2$. The prefactor is almost independent of $\Delta$, with only weak dependence near isotropy $d=1$, where Eq. (1) is recovered.
VI Summary and Outlook
@@
In summary, we have demonstrated for up to four channels that our proposed spin-chain multichannel Kondo model produces a boundary entropy $g$ consistent with Eq. 1. Thus, a fermionic background is not necessary for observing the multichannel Kondo effect. Therefore, the model can be experimentally realized in quantum computers and simulators [69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 32] where the scaling of impurity total spin, Eq. (9), might be a straightforward first step as it requires only two-qubit measurement. Additionally, near the XX limit $\Delta\to 0$, we find $J_{2}^{c}\to 0$, eliminating the need for next-nearest-neighbor interactions [66].
@@
Our numerical results pose open questions especially for the anisotropic $\Delta\neq 1$ case. In particular, a correspondence between our spin chain model and the Luttinger liquid Kondo model might shed light on the approximate power law relation between $g$ and the effective Luttinger liquid parameter $1/(2d)$ [68], see Fig. 3. Generalizing to multiple channels the Kondo model in a one-channel Luttinger liquid [79, 80, 81, 82] (possibly related to quantum Brownian motion [83, 84]) might explain the observed power law.
VII Acknowledgments
@@
We have benefited from discussions with Abay Zhakenov, Alexei Tsvelik, Chao-Ming Jian, Elio KÜnig, Mengxing Ye, Oleg Starykh, Pradip Kattel, Rui Wang, Tigran Sedrakyan, Yao Zhou and Yicheng Tang. Numerical calculations were performed by using the TeNPy Library [60]. This work was supported by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Quantum Science Center. J.G. would like to acknowledge the Office of Undergraduate Research at Purdue University for financial support. He was also financially supported through the Joel Spira Summer Research Award.
@@ 1 Department of Physics and Astronomy, Purdue University, West Lafayette, Indiana 47907, USA
These supplementary materials contain details about $J_{2}(\Delta)$ and $s_{1}(\Delta)$ in Sec. S1 and the 3CK $g(\Delta)$ values for all $J_{2}=0$ in Sec. S2.

### v2

@@ V Total Impurity Spin
Another signature of MCK physics in our model is exhibited by the total spin of the impurity and image impurity. We first find that the total spin of the whole system is spin 0 for all $0<J_{K}^{\prime}=J_{K}\leq 1$ and $M=1,2,3$. This singlet formed by the impurity spin and the rest of the spins is similar to the electronic multichannel Kondo model where the impurity spin was connected to entanglement negativity [77]. The negativity was also numerically calculated for the spin-chain 1CK case [78]. We calculate the total spin of the impurity and the image impurity, or equivalently the total spin of the $M\times L$ bulk spins because the whole system has spin 0. Due to the $\text{SU}(2)$ symmetry, it is sufficient to calculate the $z$-component,
$\langle(\mathrm{S}^{z}_{\text{tot}})^{2}\rangle=\langle(\mathrm{S}_{0}^{z}+\mathrm{S}_{0}^{{}^{\prime}z})^{2}\rangle=\frac{1}{2}+2\langle\mathrm{S}_{0}^{z}\mathrm{S}_{0}^{{}^{\prime}z}\rangle\,$, (10)
where $\langle\mathrm{S}_{0}^{z}\mathrm{S}_{0}^{z}\rangle=\langle\mathrm{S}_{0}^{{}^{\prime}z}\mathrm{S}_{0}^{{}^{\prime}z}\rangle=1/4$ by $\text{SU}(2)$ symmetry.
@@
We find that the total impurity spin Eq. (10) depends on the system size and Kondo coupling through the ratio $L/\xi_{K}$. The finite-size scaling for $\langle(\mathrm{S}^{z}_{\text{tot}})^{2}\rangle$ is shown in Fig. 2(b). In the $L/\xi_{K}\to\infty$ limit, $\langle(\mathrm{S}^{z}_{\text{tot}})^{2}\rangle\to 1/2$, which corresponds to uncorrelated impurity and image impurity. In the bulk, we get exactly the same value for the $z$-component total spins, which suggests that two independent spin-1/2 Kondo clouds are formed in the bulk. Because $\xi_{K}$ is the Kondo-cloud size, we can think of it the following way: One of the Kondo clouds is for the impurity, the other is for the image-impurity and they are decoupled from each other because their sizes $\xi_{K}$ are very small compared to the chain length $L$. In this limit, we find $\langle(\mathrm{S}^{z}_{\text{tot}})^{2}\rangle$ to deviate from 1/2 in a Fermi liquid power law $1/(L/\xi_{K})^{2}$ for $M=1$ but non-Fermi liquid power laws $1/(L/\xi_{K})$ for $M=2$ and $1/(L/\xi_{K})^{4/5}$ for $M=3$, see Fig. 2(b). These finite-size corrections can be understood to arise in second-order perturbation theory due to the least irrelevant operator with scaling dimension $2$ for $M=1$ and $1+2/(M+2)$ for $M\geq 2$ [15, 16]. A similar scaling behavior can be found in Ref. [77] for the negativity of the impurity and the bulk electrons. In the opposite limit $L/\xi_{K}\to 0$, the correlation between the impurity and the image-impurity mediated by their Kondo clouds becomes strong. Eventually, the impurity and the image-impurity form a singlet and so do $M\times L$ bulk spins.
VI Effect of Anisotropy
@@
Figure 3: (a) The effective ground state degeneracy $g=\exp(S_{\text{imp}})$ versus XXZ anisotropy parameter $d$ or $\Delta$ for $M=1,2,3,4$ and $L=50,50,50,30$ respectively. The power law $g\sim d^{(M-1)/2}$ is observed where the fit lines are for the $\Delta\leq 0$. The fit does not work well near $\Delta=1$, where it yields $g\sim\sqrt{M}$ instead of Eq. (1). (b) The decay exponent $p$ of the correlation $\langle S_{0}^{z}S_{0}^{\prime z}\rangle$ ($\sim 1/L^{p}$) and $\langle S_{0}^{x}S_{0}^{\prime x}\rangle$ (inset) versus $d$ or $\Delta$. The errors are given by the largest and smallest decay exponent fitted using two nearby even $L$ values from $92$ to $108$ for $M=1$, $42$ to $58$ for $M=2$ and $26$ to $42$ for $M=3$.
@@
We use DMRG simulations to find $g(M,\Delta)$ for $M=1,\dots,4$, shown in Fig. 3(a). We find that $g(M,\Delta)\sim d(\Delta)^{\alpha}$ closely follows a power law with the exponent $\alpha\approx(M-1)/2$. The prefactor is almost independent of $\Delta$, with only weak dependence near isotropy $d=1$, where Eq. (1) is recovered.
The power-law behaviors of the $S_{\text{imp}}$ in Fig. 2(a) are affected by the XXZ anisotropy [79]. However, in order to better demonstrate the interplay of multichannel Kondo effects and the Luttinger liquid impurity physics [47, 48, 49, 50, 51, 52, 53, 54], we show in Fig. 3(b) that the decay exponent $p$ of $\langle S_{0}^{z}S_{0}^{\prime z}\rangle\sim 1/L^{p}$ [equivalent to $\langle S_{0}^{x}S_{0}^{\prime x}\rangle$ at $\Delta=1$ due to SU(2) symmetry, see the inset of Fig. 3(b)] is almost a constant $p(\Delta)=2$ for $M=1$ and $p$ transitions from $p=2$ to $p=4/(M+2)$ as $d$ goes from 0 to $1$ for both the $M=2,3$ cases. The $M=2$ case is exactly the periodic XXZ chain and the dominant decay of $\langle S^{z}S^{\prime z}\rangle$ is $1/L^{2}$ for $\Delta<0$ and $1/L^{1/d}$ for $\Delta>0$ [82]. Currently, there is no analytical analysis on the $M=3$ case but a similar transition also occurs near $\Delta=0$, see Fig. 3(a)-(b). The decay exponent of $\langle S_{0}^{x}S_{0}^{\prime x}\rangle$ is closely $p(\Delta)=2d(\Delta)$ for $M=1$ and $p(M,\Delta)=[4/(M+2)]d(\Delta)$ for both $M=2$ [80] and $M=3$. As illustrated above, the exponent $4/(M+2)$ at $\Delta=1$ is determined by the scaling dimension of the least irrelevant operator $1+2/(M+2)$ in multichannel Kondo physics [15, 16]. With anisotropy $\Delta$, all the $M=1,2,3$ exponents are reduced by a factor $d(\Delta)$.
VII Summary and Outlook
@@
In summary, we have demonstrated for up to four channels that our proposed spin-chain multichannel Kondo model produces a boundary entropy $g$ consistent with Eq. 1. Thus, a fermionic background is not necessary for observing the multichannel Kondo effect. Therefore, the model can be experimentally realized in quantum computers and simulators [83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 35] where the scaling of impurity total spin, Eq. (10), might be a straightforward first step as it requires only two-qubit measurement. Note that our discussion is limited to $J_{K}\leq 1$ because the model with $J_{K}>1$ may enter non-Kondo phases [93].
@@
Our numerical results pose open questions especially for the anisotropic $\Delta\neq 1$ case. Generalizing to multiple channels the Kondo model in a one-channel Luttinger liquid [94, 95, 96, 97, 48, 49] (possibly related to quantum Brownian motion [98, 99]) might explain the observations in Fig. 3. Other future directions include exploring possible relation to two-impurity spin-chain Kondo models, extending the model to other quantum impurity models such as non-Hermitian spin-chain multichannel Kondo models [100], and studying finite-temperature and dynamical (finite-frequency) response functions.
VIII Acknowledgments
@@
We have benefited from discussions with A. Zhakenov, A. Tsvelik, C.-M. Jian, E. König, M. Ye, O. Starykh, P. Kattel, R. Wang, T. Sedrakyan, Y. Zhou and Y. Tang. Numerical calculations were performed by using the TeNPy Library [72]. This work was supported by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Quantum Science Center. J.G. would like to acknowledge the Office of Undergraduate Research at Purdue University for financial support. He was also financially supported through the Joel Spira Summer Research Award.
@@ 1 Department of Physics and Astronomy, Purdue University, West Lafayette, Indiana 47907, USA
These supplementary materials contain details about $J_{2}(\Delta)$ and $s_{1}(\Delta)$ in Sec. S1 and the 3CK $g(\Delta)$ values for all $J_{2}=0$ in Sec. S2. Finally, we investigate the influence of anisotropy $\Delta$ on the impurity entropy scaling in Sec. S3, and in Sec. S4 we evaluate spin-spin correlation functions between the impurity and bulk spins.
@@
S3 The $S_{\text{imp}}-\ln(g)$ power law v.s. anisotropy $\Delta$
We calculated the $S_{\text{imp}}(x)$ with $x=L/4$ by taking $\Delta\in[-0.9,-0.8,\dots,1]$ and $L\in[44,48,52,56]$ for $M=1,2$ and $L\in[28,32,36,40]$ for $M=3$. By fitting the data we obtained, we get the exponent $p$ in $S_{\text{imp}}-\ln[g(M,\Delta)]\sim 1/L^{p}$ where $g(1,\Delta)=1$, $g(2,\Delta)=\sqrt{2}d^{1/2}$ and $g(3,\Delta)=\frac{\sqrt{5}+1}{2}d(\Delta)$ with $d(\Delta)=1-\arccos(\Delta)/\pi$. The $\Delta=1$ results for $M=1,3$ in the following Fig. S3 match the expected exponent $1$ for $M=1$ and $4/5$ for $M=3$. The $M=2,\Delta=1$ behavior is $\ln(L)/L$, not the power law $1/L$ and thus a lower effective power-law scaling $\sim 0.6$ is observed. We did not find any simple dependence on $d$ and $\Delta$ for the exponent $p$, not like what we found in Fig. 3(b) for the correlation functions.
Figure S3: The scaling exponent $p$ for $S_{\text{imp}}-\ln(g)\sim 1/L^{p}$ versus anisotropy parameter $d$ or $\Delta$ with $J_{2}=J_{2}^{c}(\Delta)$ for $M=1,2,3$.
S4 The $\langle S^{x}_{0}S^{k=1,x}_{r}\rangle$ power law for different anisotropy parameters
In order to show the Luttinger liquid physics, we also calculate the correlation functions between the impurity spin $S_{0}$ and the spin in the $k$th channel and at position $r=1,\dots,L,L+1$ where $r=L+1$ denotes image-impurity. For example, the x-component can be written as $\langle S^{x}_{0}S^{k=1,x}_{r}\rangle=\langle S^{x}_{0}S^{k=1,x}_{r}\rangle_{U}+(-1)^{r}\langle S^{x}_{0}S^{k=1,x}_{r}\rangle_{A}$ as a function of $r$ and the alternating part $\langle S^{x}_{0}S^{k=1,x}_{r}\rangle_{A}$ is shown in Fig. S4. The $M=2$ case (periodic XXZ chain) matches with the prediction in Ref. [80].
Figure S4: A log-log plot of $\langle S^{x}_{0}S^{k=1,x}_{r}\rangle_{A}$ v.s. $r$ for $M=1,2,3$ (red, blue, green) and $L=98,58,42$ respectively. The Luttinger-liquid scaling behavior is clearly seen from the power-law lines with exponents $(3/2)d,d,(4/5)d$ for $M=1,2,3$ and the Luttinger parameter is $K=1/(2d)$.